OpenStudy (anonymous):

SOLVE FOR t : 9/10t = 90

OpenStudy (anonymous):

9=90x10t 9=900t 900/9=t 100=t

OpenStudy (anonymous):

\[\frac{9}{10t} = 90\] or \[\frac{9}{10}t = 90\] ?